OpenStudy (meilendurcer):

Determine which system below will produce infinitely many solutions. My answer is the last one can someone check these? -4x + 6y = 20 4x - 3y = -2 -2x + 4y = 18 4x - 8y = -36 3x - 2y = 5 -6x - 4y = -10 5x - 3y = 12 15x - 9y = 24 The art museum had a total of 224 visitors on Tuesday. Visitors older than 18 paid $12 for admission. Visitors 18 years or younger paid $8 for admission. The museum collected a total of $2,520. How many visitors were 18 or younger? 42 56 182 194 my answer is 194

OpenStudy (anonymous):

Sorry the answer for 1st one is -2x + 4y = 18 4x - 8y = -36

OpenStudy (meilendurcer):

ahh ok i was going back and forth on those answers

OpenStudy (anonymous):

That's incorrect Salmon She had it right.

OpenStudy (anonymous):

2nd one equations are x+y=224 12x+8y=2520

OpenStudy (anonymous):

Oops nevermind yes the 2nd one has infinite solutions The last has no solutions.

OpenStudy (meilendurcer):

do i put 12 and 8 into the equations?

OpenStudy (anonymous):

no...Rearrange equation one as x=224-y now substitute it in equation two....

OpenStudy (meilendurcer):

can you help i don't understand about substituting and all

OpenStudy (anonymous):

after substituting equation two becomes 2688-12y+8y=2520 4y=168 y=42 Got it???

OpenStudy (meilendurcer):

yes i think i do.

OpenStudy (meilendurcer):

thanks